Global Manufacturing Company purchased new equipment on April 25, 2021, at a cost of $80000. 

Useful life of this equipment was estimated at 4 years, with an estimated residual value of $5000. For 

income tax purpose, this equipment is classified as “5-years property”.

Instructions: Compute the annual depreciation expense for each year until this equipment becomes 

fully depreciated under each of depreciation method listed below.

i) Straight-line, with depreciation for fractional years rounded to the nearest whole month.

ii) 200%- declining-balance, with the half-year convention


1
Expert's answer
2021-08-16T14:47:21-0400

i)

2021


"\frac{80000-5000}{44}\times8=13636"

2022-2024


"\frac{80000-5000}{44}\times12=20455"

ii)

2021


"\frac{80000-2*5000}{7}=10000"

2022


"\frac{80000-2*10000}{6}=10000"

"\frac{80000-2*10000}{5}=12000"

2023


"\frac{80000-2*12000}{4}=14000"

"\frac{80000-2*14000}{3}=17333"

2024


"\frac{80000-2*17333}{2}=22667"

"\frac{80000-2*22667}{1}=34666"